OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x - 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (anonymous):

I just need to know the steps on how to do this. Can someone guide me? @skullpatrol @ranga @BTaylor

OpenStudy (ranga):

Is this a calculus question? If it is a calculus question then you can find the derivative dT/dx, set it to zero and solve for x. If this is a pre-calculus question, then you can answer is it as follows: The function T = x^3 has a turning point at x = 0. We can determine this simply by looking at the graph T = x^3. Therefore, (0,0) is the turning point for the graph T = x^3. The function T(x) = (x - 4)^3 + 6 will be the graph T = x^3 shifted to the right by 4 units and shifted up by 6 units. Therefore, the turning point will also shift from (0,0) to 4 units to the right and 6 units up. The turning point will be (4,6) for the function T(x) = (x - 4)^3 + 6.

OpenStudy (anonymous):

It actually is an Algebra 2 question. Will the above still work? @ranga

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Okay thanks!